OpenStudy (anonymous):

.6x-.1=.5x+.2 to get rid of the decimal i can multiply by 10 right ? but do i multiply everysingle term?

OpenStudy (chaise):

Yes. Every single term.

OpenStudy (anonymous):

think of it this way you know \[2+11=9+4\] and and if you want to multiply by ten and still get a true equality it must be \[20+110=90+40\]

OpenStudy (anonymous):

you can't multiply just one part by ten and not all of it